2 cd's and 3 dvd's cost $70.97. 3 cd's and 2 dvd's cost $66.48. What is the cost of each cd and DVD?
Answer by TimothyLamb(4379)

x = dollar price of one CD
y = dollar price of one DVD
---
2x + 3y = 70.97
3x + 2y = 66.48
---
put the system of linear equations into standard form
---
2x + 3y = 70.97
3x + 2y = 66.48
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 11.50
y = 15.99
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---